# Composing Global Optimizers to Reasoning Tasks via Algebraic Objects in Neural Nets

### v3

@@ C.3 Canonical Forms
Since $r_{k_{0}k_{0}k_{0}}({\bm{z}})=\sum_{j}a_{k_{0}j}b_{k_{0}j}c_{k_{0}j}\neq 0$, there must exist some $j$ so that $z_{ak_{0}j}z_{bk_{0}j}z_{ck_{0}j}\neq 0$, which means that $z_{ak_{0}j}\neq 0$, $z_{bk_{0}j}\neq 0$ and $z_{ck_{0}j}\neq 0$. Since the node index $j$ can be permuted, we can let node $j$ be the first node 00 and let $y_{pk0}=z_{pkj}$ and $z^{\prime}_{pkj^{\prime}}=z_{pkj^{\prime}}z_{pkj}^{-1}$ for $p\in\{a,b,c\}$ and $k=\pm k_{0}$, then ${\bm{z}}^{\prime}$ is canonical at $k_{0}$ and $\mathrm{ord}({\bm{y}})=1$. Finally, by ring homomorphism, since
@@
$(\theta_{a},\theta_{b},\theta_{c})$ =(±$=(\pm\pi/2,0,0)\mod 2\pi$ (78)
@@
where $\Re(\cdot)$ and $\Im(\cdot)$ are real and imaginary parts of a complex vector. Since Eqn. 88 holds in $\mathbb{R}^{3}$, it must be the following cases: either $\Re({\bm{a}}\circ\bar{\bm{b}})$ is co-linear with $\Im({\bm{a}}\circ\bar{\bm{b}})$, or $\Re({\bm{c}})$ is co-linear with $\Im({\bm{c}})$.
@@
Remarks. Note that this conclusion does not contradict with the constructed solution ${\bm{z}}_{\mathrm{syn},\alpha\beta}$ in Eqn. 11 in which $\alpha$ and $\beta$ are allowed to be any complex number with magnitude $1$. This is because ${\bm{z}}_{\mathrm{syn},\alpha\beta}$ does not satisfy all the constraints in $R_{\mathrm{c}}$ (but ${\bm{z}}_{\mathrm{syn},\alpha\beta}*{\bm{z}}_{\nu=\mathrm{i}}$ will) unless $\alpha$ and $\beta$ are real and thus ±$\pm 1$.
Appendix D Gradient Dynamics (Proof)
See 5
@@ Appendix D Gradient Dynamics (Proof)
$-\overline{\nabla_{\bm{r}}\ell}^{*}\dot{\bm{r}}=\overline{\nabla_{\bm{r}}\ell}^{*}JJ^{*}\overline{\nabla_{\bm{r}}\ell}=\|J^{*}\overline{\nabla_{\bm{r}}\ell}\|_{2}^{2}\geq 0$ (100)
See 6
@@
$\hat{\bm{y}}(0;\lambda)={\bm{e}},\quad\quad\hat{\bm{y}}(1;\lambda)={\bm{y}},\quad\quad r_{kkk}(\hat{\bm{y}}(t;0))\leq 1,\quad\quad r_{kkk}(\hat{\bm{y}}(t;1))\leq 1$ (102)
@@
Let ~$\tilde{\ell}:=\sum_{k}\nabla\tilde{\ell}_{k}$. Let’s compute the dynamics of MPs following Theorem 5: $\dot{\bm{r}}=-JJ^{*}\overline{\nabla_{\bm{r}}\tilde{\ell}}$.
@@
~$\frac{\partial\tilde{\ell}}{\partial r_{k_{1}k_{2}k}}=\sum_{k}\frac{\partial\tilde{\ell}_{k}}{\partial r_{k_{1}k_{2}k}}=-2\mathbb{I}(k_{1}=k_{2}=k)+2\overline{r_{k_{1}k_{2}k}}$ (108)
@@
where $\kappa>0$ is the diagonal element of $JJ^{*}$ and $\alpha$ is a coefficient that characterizes the relative strength of two negative gradient $-\overline{\nabla_{r_{kkk}}\ell}=r_{0}-r_{kkk}$ and $-\overline{\nabla_{r_{k_{1}k_{2}k}}\ell}=-r_{k_{1}k_{2}k}$, and $c_{0}$ is the gradient terms caused by asymmetry and/or other frequencies. This yields a second-order ODE that has complex roots in the characteristic function when $c_{0}>0$.
Appendix E Extending CoGO to Group Action Prediction
While in this work we mainly focus on Abelian group, CoGO can be extended to more general group action prediction: given a group element $g\in G$ and the current state $x\in\mathcal{X}$, the goal is to predict $gx\in X$, i.e., the next state after action $g$. Such tasks include modular addition/multiplication in which the group acts on itself (i.e., $\mathcal{X}=G$), and also includes the transition function in reinforcement learning (Sutton, 2018) and world modeling (Garrido et al., 2024), in which an action changes the current state to a new one.
@@ Appendix E Extending CoGO to Group Action Prediction
Moreover, we can prove that for any group element $g\in G$, there exists $x=\iota_{0}(g)\in\mathcal{X}$ so that for any state $x^{\prime}$, the group action $gx^{\prime}$ is the same as the state multiplication $x^{\prime}\cdot x$. Therefore, for group action prediction tasks, we have (note the difference compared to Eqn. 12):
@@
Appendix F Detailed explanation of Sec. E
@@ F.1 Transitive Case
There is a bijection between $\{\iota_{0}^{-1}(x)\}_{x\in\mathcal{X}}$ and co-sets $[G:G_{x_{0}}]$ of group stabilizer $G_{x_{0}}:=\{g\in G|gx_{0}=x_{0}\}$, which is a subgroup of $G$ fixing $x_{0}$.
@@
F.2 General group action
@@
Appendix G Proofs for the content in Appendix
See 7
@@ Appendix G Proofs for the content in Appendix
So any $g$ is a bijective mapping on $\mathcal{X}$. Since every element of $R_{g}$ is either 00 or $1$, $R_{g}$ is a permutation matrix. ∎
See 8
@@
Appendix H Additional Experiments
Algorithm to extract factorization from gradient descent solutions. Given the solutions obtained by gradient descent using Adam optimizer, we first compute the corresponding ${\bm{z}}$ via the Fourier transform (that is, Eqn. 12). Here ${\bm{z}}=[z_{pkj}]$ is a $3$-by-$d$-by-$q$ tensor. Here $d=|G|$ and $q$ is the number of hidden nodes in the 2-layer neural networks.
@@ Appendix H Additional Experiments
Suppose we now get ${\bm{z}}^{(k)}$ for frequency $k$, which is a $3$-by-$6$ (and thus an order-6) solution. Then we enumerate all possible permutation of $6$ hidden nodes ($6!=720$ possibilities) to find one permutation $\tau$ so that $\|z_{pk\tau(\cdot)}-z^{(1)}_{pk\cdot}\otimes z^{(2)}_{pk\cdot}\|$ is minimized, following ring multiplication defined in Def. 5. Note that for each permutation, we also need to consider whether $\tilde{\bm{1}}:=[-1,-1,1]$ can be applied to each hidden node $j$ ($\tilde{\bm{1}}$ is also defined in Tbl. 1). This is because both ${\bm{z}}_{1}+{\bm{z}}_{2}$ and ${\bm{z}}_{1}+\tilde{\bm{1}}*{\bm{z}}_{2}$ have exactly the same values on all monomial potentials (MPs) we consider, due to the fact that $r(\tilde{\bm{1}})=1$ for any $r\in R_{\mathrm{g}}\cup R_{\mathrm{c}}\cup R_{\mathrm{n}}\cup R_{*}$. Therefore we call $\tilde{\bm{1}}$ “pseudo-1”.
@@
Once we find such ${\bm{z}}_{1}$ and ${\bm{z}}_{2}$, we convert them into their canonical forms $\tilde{\bm{z}}_{1}$ and $\tilde{\bm{z}}_{2}$ (Def. 8) to eliminate any possible multiplicative term ${\bm{y}}$ so that ${\bm{z}}_{1}={\bm{y}}*\tilde{\bm{z}}_{1}$. We then compare the canonical forms (up to complex conjugate) with various order-3 and order-2 partial solutions constructed by CoGO, as detailed in Sec. 5. If their distance is below a certain threshold (e.g., $<10\%$ of the norm after normalizing both $\hat{\bm{z}}_{1}$ and $\hat{\bm{z}}_{2}$), then a match is detected.

### v4

@@ D.3 Canonical Forms
Since $r_{k_{0}k_{0}k_{0}}({\bm{z}})=\sum_{j}a_{k_{0}j}b_{k_{0}j}c_{k_{0}j}\neq 0$, there must exist some $j$ so that $z_{ak_{0}j}z_{bk_{0}j}z_{ck_{0}j}\neq 0$, which means that $z_{ak_{0}j}\neq 0$, $z_{bk_{0}j}\neq 0$ and $z_{ck_{0}j}\neq 0$. Since the node index $j$ can be permuted, we can let node $j$ be the first node 0 and let $y_{pk0}=z_{pkj}$ and $z^{\prime}_{pkj^{\prime}}=z_{pkj^{\prime}}z_{pkj}^{-1}$ for $p\in\{a,b,c\}$ and $k=\pm k_{0}$, then $\bm{z}$$^{\prime}$ is canonical at $k_{0}$ and $\mathrm{ord}({\bm{y}})=1$. Finally, by ring homomorphism, since
@@
$(\theta_{a},\theta_{b},\theta_{c})$ $=(\pm\pi/2,0,0)\mod 2\pi$ (77)
@@
where $\Re(\cdot)$ and $\Im(\cdot)$ are real and imaginary parts of a complex vector. Since Eqn. 87 holds in $\mathbb{R}^{3}$, it must be the following cases: either $\Re({\bm{a}}\circ\bar{\bm{b}})$ is co-linear with $\Im({\bm{a}}\circ\bar{\bm{b}})$, or $\Re({\bm{c}})$ is co-linear with $\Im({\bm{c}})$.
@@
Remarks. Note that this conclusion does not contradict with the constructed solution ${\bm{z}}_{\mathrm{syn},\alpha\beta}$ in Eqn. 68 in which $\alpha$ and $\beta$ are allowed to be any complex number with magnitude $1$. This is because ${\bm{z}}_{\mathrm{syn},\alpha\beta}$ does not satisfy all the constraints in $R_{\mathrm{c}}$ (but ${\bm{z}}_{\mathrm{syn},\alpha\beta}*{\bm{z}}_{\nu=\mathrm{i}}$ will) unless $\alpha$ and $\beta$ are real and thus $\pm 1$.
Appendix E Gradient Dynamics (Proof)
[…]
Let ${\bm{r}}=[r_{k_{1}k_{2}k},r_{pk_{1}k_{2}k}]\in\mathbb{C}^{4d^{3}}$ be a vector of all MPs, and $J:=\frac{\partial{\bm{r}}}{\partial{\bm{z}}}\frac{\partial{\bm{z}}}{\partial\mathcal{W}}$ be the Jacobian matrix of the mapping ${\bm{r}}={\bm{r}}({\bm{z}}(\mathcal{W}))$ in which $\mathcal{W}$ is the collection of original weights. Note that when we take derivatives with respect to $r$ and apply chain rules, we treat $r$ and its complex conjugate (e.g., $r_{kkk}$ and $r_{-k,-k,-k}=\bar{r}_{kkk}$) as independent variables. Since we run the gradient descent on $\mathcal{W}$, will such (indirect) optimization leads to a descent of $\bm{r}$} towards the desired targets (Lemma 1)? The following lemma confirms that:
Lemma 7 (Dynamics of MPs).
The dynamics of MPs satisfies $\dot{\bm{r}}=-JJ^{*}\overline{\nabla_{\bm{r}}\ell}$, which has positive inner product with the negative gradient direction $\overline{\nabla_{\bm{r}}\ell}$.
@@ Appendix E Gradient Dynamics (Proof)
$\overline{\nabla_{\bm{r}}\ell}^{*}\dot{\bm{r}}=\overline{\nabla_{\bm{r}}\ell}^{*}JJ^{*}\overline{\nabla_{\bm{r}}\ell}=\|J^{*}\overline{\nabla_{\bm{r}}\ell}\|_{2}^{2}\geq 0$ (99)
See 5
@@
$\hat{\bm{y}}(0;\lambda)={\bm{e}},\quad\quad\hat{\bm{y}}(1;\lambda)={\bm{y}},\quad\quad r_{kkk}(\hat{\bm{y}}(t;0))\leq 1,\quad\quad r_{kkk}(\hat{\bm{y}}(t;1))\geq 1$ (101)
@@
Let $\tilde{\ell}:=\sum_{k}\nabla\tilde{\ell}_{k}$. Let’s compute the dynamics of MPs following Theorem 7: $\dot{\bm{r}}=-JJ^{*}\overline{\nabla_{\bm{r}}\tilde{\ell}}$.
@@
$\frac{\partial\tilde{\ell}}{\partial r_{k_{1}k_{2}k}}=\sum_{k}\frac{\partial\tilde{\ell}_{k}}{\partial r_{k_{1}k_{2}k}}=-2\mathbb{I}(k_{1}=k_{2}=k)+2\overline{r_{k_{1}k_{2}k}}$ (107)
@@
where $\kappa>0$ is the diagonal element of $JJ^{*}$ and $\alpha$ is a coefficient that characterizes the relative strength of two negative gradient $\overline{\nabla_{r_{kkk}}\ell}=r_{0}-r_{kkk}$ and $\overline{\nabla_{r_{k_{1}k_{2}k}}\ell}=-r_{k_{1}k_{2}k}$, and $c_{0}$ is the gradient terms caused by asymmetry and/or other frequencies. This yields a second-order ODE that has complex roots in the characteristic function when $c_{0}>0$.
Appendix F Extending CoGS to Group Action Prediction
While in this work we mainly focus on Abelian group, CoGS can be extended to more general group action prediction: given a group element $g\in G$ and the current state $x\in\mathcal{X}$, the goal is to predict $gx\in X$, i.e., the next state after action $g$. Such tasks include modular addition/multiplication in which the group acts on itself (i.e., $\mathcal{X}=G$), and also includes the transition function in reinforcement learning (Sutton, 2018) and world modeling (Garrido et al., 2024), in which an action changes the current state to a new one.
@@ Appendix F Extending CoGS to Group Action Prediction
Moreover, we can prove that for any group element $g\in G$, there exists $x=\iota_{0}(g)\in\mathcal{X}$ so that for any state $x^{\prime}$, the group action $gx^{\prime}$ is the same as the state multiplication $x^{\prime}\cdot x$. Therefore, for group action prediction tasks, we have (note the difference compared to Eqn. 11):
@@
Appendix G Detailed explanation of Sec. F
@@ G.1 Transitive Case
There is a bijection between $\{\iota_{0}^{-1}(x)\}_{x\in\mathcal{X}}$ and co-sets $G:G_{x_{0}}]$ of group stabilizer $G_{x_{0}}:=\{g\in G|gx_{0}=x_{0}\}$, which is a subgroup of $G$ fixing $x_{0}$.
@@
G.2 General group action
@@
Appendix H Proofs for the content in Appendix
See 8
@@ Appendix H Proofs for the content in Appendix
So any $g$ is a bijective mapping on $\mathcal{X}$. Since every element of $R_{g}$ is either 0 or $1$, $R_{g}$ is a permutation matrix. ∎
See 9
@@
Appendix I Additional Experiments
Algorithm to extract factorization from gradient descent solutions. Given the solutions obtained by gradient descent using Adam optimizer, we first compute the corresponding $\bm{z}$} via the Fourier transform (that is, Eqn. 11). Here ${\bm{z}}=[z_{pkj}]$ is a 33-by-$d$-by-$q$ tensor. Here $d=|G|$ and $q$ is the number of hidden nodes in the 2-layer neural networks.
@@ Appendix I Additional Experiments
Suppose we now get ${\bm{z}}^{(k)}$ for frequency $k$, which is a 33-by-66 (and thus an order-6) solution. Then we enumerate all possible permutation of $6$ hidden nodes ($6!=720$ possibilities) to find one permutation $\tau$ so that $\|z_{pk\tau(\cdot)}-z^{(1)}_{pk\cdot}\otimes z^{(2)}_{pk\cdot}\|$ is minimized, following ring multiplication defined in Def. 5. Note that for each permutation, we also need to consider whether $\tilde{\bm{1}}:=[-1,-1,1]$ can be applied to each hidden node $j$ ($\tilde{\bm{1}}$ is also defined in Tbl. 1). This is because both ${\bm{z}}_{1}+{\bm{z}}_{2}$ and ${\bm{z}}_{1}+\tilde{\bm{1}}*{\bm{z}}_{2}$ have exactly the same values on all sum potentials (SPs) we consider, due to the fact that $r(\tilde{\bm{1}})=1$ for any $r\in R_{\mathrm{g}}\cup R_{\mathrm{c}}\cup R_{\mathrm{n}}\cup R_{*}$. Therefore we call $\tilde{\bm{1}}$ “pseudo-1”.
@@
Once we find such ${\bm{z}}_{1}$ and ${\bm{z}}_{2}$, we convert them into their canonical forms $\tilde{\bm{z}}_{1}$ and $\tilde{\bm{z}}_{2}$ (Def. 8) to eliminate any possible multiplicative term $\bm{y}$} so that ${\bm{z}}_{1}={\bm{y}}*\tilde{\bm{z}}_{1}$. We then compare the canonical forms (up to complex conjugate) with various order-3 and order-2 partial solutions constructed by CoGS, as detailed in Sec. 5. If their distance is below a certain threshold (e.g., $10\%$ of the norm after normalizing both $\hat{\bm{z}}_{1}$ and $\hat{\bm{z}}_{2}$), then a match is detected.
@@
Figure 12: Distribution of free parameters ($\xi$, $\nu$, $\alpha$ and $\beta$, all with magnitude 11) in three kinds of gradient descent solutions identified by CoGS. While any value of these parameters makes a global solution, gradient descent dynamics has a particular preference in picking them during optimization.